Fulkerson Manufacturing wishes to maintain a sustainable growth rate of 9.25 percent a year, a debt–equity ratio of .40, and a dividend payout ratio of 32.5 percent. The ratio of total assets to sales is constant at 1.35. What profit margin must the firm achieve in order to meet its growth rate goal? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Profit margin %
Answer #1

Sustainable Growth Rate = 9.25%
Payout Ratio = 32.50%

Retention Ratio, b = 1 - Payout Ratio
Retention Ratio, b = 1 - 0.3250
Retention Ratio, b = 0.6750

Sustainable Growth Rate = [ROE * b] / [1 - ROE * b]
0.0925 = [ROE * 0.675] / [1 - ROE * 0.675]
0.0925 - 0.06244 * ROE = 0.675 * ROE
0.0925 = 0.73744 * ROE
ROE = 0.1254 or 12.54%

Equity Multiplier = 1 + Debt-Equity Ratio
Equity Multiplier = 1 + 0.40
Equity Multiplier = 1.40

Assets Turnover = Sales / Total Assets
Assets Turnover = 1 / 1.35
Assets Turnover = 0.7407

ROE = Profit Margin * Assets Turnover * Equity Multiplier
0.1254 = Profit Margin * 0.7407 * 1.40
Profit Margin = 0.1209 or 12.09%